how many liters of pure water must be combined with an 18-liter mixture that is
24% acid in order to obtain a mixture that is 15% acid? Your answer may be a
decimal number, but it must be exact.
Answer: ________________________________

Answer by mananth(16946)

--- percent ---------------- Amount
acid 24 ---------------- 18 liters
water 0 ---------------- x liters
Mixture 15 ---------------- 18 + x liters

24 * 18 + 0 x = 15 ( 18 + x )
432 + 0 x = 270 + 15 x
0 x -15 x = -432 + 270
-15 x = -162
/ -15
x = 10.8 liters water